OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 10 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

x-y = 10.........(i) 2x+y = 2 .........(ii) from (i) x = 10+y substitute in (ii) 2(10+y) + y = 2 20+2y+y = 2 3y = 2-20 3y = -18 y = -6, hope you following, now can you find x?